\begin{thebibliography}{9}
\itemsep=-2pt
\small
\bibitem{AMOT} D. Aoki, K. Minematsu, T. Okamura and T. Takagi, Efficient Word Size Modular Multiplication over Signed Integers, IEEE 29th Symposium on Computer Arithmetic (ARITH), Lyon, France, 2022, pp. 94-101.
\bibitem{BM} J. W. Bos and P. L. Montgomery. Montgomery Arithmetic from a Software Perspective. In: Bos JW, Lenstra AK, eds. Topics in Computational Number Theory Inspired by Peter L. Montgomery. Cambridge University Press; 2017:10-39. 
%\bibitem{fft} Cooley J. W., Tukey, J. W. An algorithm for the
%machine calculation of complex Fourier series,
%Math. Comput., 19(1965),  297-301.
\bibitem{dlx} G. Davida, B. Litow , and G. Xu, Fast arithmetics using Chinese Remaindering, {\sl Information Processing
Letters}, 109(2009), 660-662.
%\bibitem{Harv} Harvey D., Faster arithmetic for number-theoretic transforms, Journal of Symbolic Computation, 60(2014),113-119.
\bibitem{Huang} J. Huang, J. Zhang, H. Zhao et al., Improved Plantard arithmetic for lattice-based cryptography,  IACR Transactions on Cryptographic Hardware and Embedded Systems, 2022, 2022(4): 614-636.
\bibitem{Huang2}  J. Huang, H. Zhao, J. Zhang, et al., Yet another Improvement of Plantard Arithmetic for Faster Kyber on Low-end 32-bit IoT Devices, IEEE Transactions on Information Forensics and Security, vol. 19, pp. 3800-3813, 2024.
\bibitem{Libb} U. Libbrecht, Chinese Mathematics in the Thirteenth Century, Dover Publications, 2006.
%\bibitem{LongaNaehrig} P. Longa and M. Naehrig, Speeding up the number theoretic transform for faster
%ideal lattice-based cryptography,  In CANS 2016, volume 10052 of Lecture Notes
%in Computer Science, pages 124-139. Springer, 2016.
\bibitem{LyuSei}
V. Lyubashevsky, and G. Seiler. NTTRU: Truly Fast NTRU Using NTT. IACR Transactions on Cryptographic Hardware and Embedded Systems, 2019(3), 180-201. https://doi.org/10.13154/tches.v2019.i3.180-201
\bibitem{Mont} P. L. Montgomery, Modular multiplication without trial division, Mathematics of Computation, 44(1985), 519-521.
%\bibitem{Nuss} H. J. Nussbaumer, Fast polynomial transform algorithms for digital %convolution,  IEEE Trans.
%Acoust. Speech Signal Process., 28 (1980), 205-215.
\bibitem{Plan} T. Plantard,  Efficient word size modular arithmetic, IEEE Transactions on Emerging Topics in Computing, 2021, 9(3): 1506-1518.
\bibitem{Scott} M. Scott, A note on the implementation of the number theoretic transform.  Cryptography and Coding: 16th IMA International Conference, 2017: 247-258.
\bibitem{SV} M. Shand and J. Vuillemin. Fast implementations of RSA cryptography. In E. E. S. Jr., M. J. Irwin, and G. A. Jullien, editors, 11th Symposium on Computer Arithmetic, pages 252-259. IEEE Computer Society, 1993.
\bibitem{qin1247} J. Qin, Mathematical Treatise in Nine Sections, 1247.
%\bibitem{Qin} Jiushao Qin, Mathematical Treatise in Nine Sections, 1247.
\bibitem{Seiler} G. Seiler, Faster AVX2 optimized NTT multiplication for
Ring-LWE lattice cryptography, https://eprint.iacr.org/2018/039.
%\bibitem{ntl} V. Shoup, NTL: a library for doing number theory (Version5.5.2), http://www.shoup.net/ntl/.
%\bibitem{vzGG} J. von zur Gathen,  J. Gerhard, Modern Computer Algebra (3rd ed.), Cambridge Press, 2013
\bibitem{Xu} G. Xu, A Remark on Fourier Transform, https://arxiv.org/abs/1807.05829, 2018.
\bibitem{YJX} Y. Yang, Y. Jia and G. Xu. On Modular Algorithms and Butterfly Operations in Number Theoretic Transform, https://arxiv.org/html/2402.00675v2, 2024.
\end{thebibliography}

\documentclass{article}
\usepackage{amssymb,amsmath,amsthm,proof}
\usepackage{graphicx}
\usepackage{gensymb}

\usepackage{multicol}
\usepackage{multirow}

\newtheorem{thm}{Theorem}[section]

\newtheorem{example}{Example}[section]

\usepackage[section]{algorithm}
\usepackage{algpseudocode,algorithmicx}

\newcommand*\Let[2]{\State #1 $\gets$ #2}

\title{
Chinese Remainder Theorem Approach to Montgomery-Type Algorithms}
\date{}
\author
{Guangwu Xu\thanks{SCST, Shandong University, China,  e-mail: {\tt gxu4sdq@sdu.edu.cn}(Corresponding author)},
	Yiran Jia\thanks{SCST, Shandong University, China,  e-mail: {\tt kawyon@mail.sdu.edu.cn}.}, Yanze Yang\thanks{SCST, Shandong University, China, e-mail: {\tt yyannze99@mail.sdu.edu.cn}.}.\\
}

\begin{document}


\baselineskip18pt


\maketitle

\begin{abstract}
This paper explores the ability of the Chinese Remainder Theorem formalism to model Montgomery-type algorithms.
A derivation of CRT based on Qin's Identity gives Montgomery reduction algorithm immediately. This establishes a unified framework to treat modular reduction algorithms  of  Montgomery-type. Several recent notable variants of Montgomery algorithm are analyzed, validation of these methods are performed within the framework. Problems in some erroneous design of
reduction algorithms  of  Montgomery-type in the literature are detected and counter examples are easily generated by using the CRT formulation.

{\bf Key words:} Modular arithmetic, Montgomery algorithm, Chinese remainder theorem.

\end{abstract}

\section{Introduction}
Modular arithmetic is a basic tool in modern computation. Its efficiency has been   critical to applications in cryptography and communication. Modular multiplication is of the most concern with respect to the cost of computation.
In 1985, Montgomery invented a fast method for modular multiplication in his seminal paper \cite{Mont}. This ingenious design makes a very clever use of the fact that integer division (and modulation) by a power of $2$ is trivial.
Given an odd integer $N>0$, one chooses an $R>N$ which is a power of $2$.  Montgomery first proposes a representation of the ordinary residues as
\[
\begin{array}{lrl}
\Phi:& \{0,1,2,\cdots, N-1 \}&\rightarrow \{0,1,2,\cdots, N-1 \}\\
&x &\mapsto xR \pmod N ,
\end{array}
\]
and performs arithmetic on the representations. This map preserves addition but not multiplication. In essence, Montgomery reduction ``pulls back'' a multiplication result to the form of representation. More specifically, with  exactly the same formulation as in  \cite{Mont}, Montgomery's reduction algorithm  REDC computes $TR^{-1}\pmod N$ for
$0\le T<RN$ and  $N' = R-(N^{-1} \mod R)$:

{\bf function}   REDC($T$) \\
${}\quad (1)\quad\quad m\gets (T \mod R) N' \mod R$\\
${}\quad (2)\quad\quad t\gets \frac{T +mN}R$\\
${}\quad (3)\quad\quad${\bf if } $t\ge N$ {\bf then return } $t-N$ {\bf else return } $t \qquad \blacksquare$

This function speeds the calculation by using only two multiplication and eliminating the ordinary division.

For a technical motivation of Montgomery algorithm, the literature regards it as a generalization of Hensel's odd division for 2-adic numbers \cite{BM,SV}. The main theme of this paper is that Montgomery algorithm can be modeled by the Chinese Remainder Theorem (CRT), so that the number theoretical nature and computational characteristic can be seen more transparently.

Recently, there has been a huge demand of using number theoretic transform (NTT) in
systems for post-quantum cryptography and fully-homomorphism encryptions, several modular reduction algorithms of Montgomery type have been proposed for such specific applications, see, for example, \cite{AMOT,Huang,Huang2,Plan,Seiler}. Like Montgomery algorithm, these variants can be naturally  treated by CRT as well.

Our CRT approach of modular reduction algorithms of Montgomery type is well tested in several ways. We setup a CRT framework whose formulation perfectly matches 
the expression of Montgomery reduction algorithm. 
Derivation and proofs of correctness for other
methods of Montgomery type
can be obtained within  CRT framework in a unified manner. Moreover, problems in some erroneous design in the literature can be easily detected, and counter examples can be easily generated, both are done with CRT formulation. 
To the best of our knowledge, this is the first formal and detailed description of CRT approach to algorithms of Montgomery type in paper, even though one of the authors had talked this idea long time ago\footnote{The idea has been well-spread as the first author discussed this approach in several public seminar talks since 2021. For example, one of which was given at the IIE of Chinese Academy of Sciences on November 25, 2021.}.

This paper is revision of our early version \cite{YJX} with focus on the single
topic of modulo
operation, without touching on the NTT applications. 

In the next section, we will first present a derivation of CRT based on Qin's Identity, a formula obtained from the ideas of an ancient book by Jiushao Qin \cite{qin1247}. By a simple manipulation, this gives Montgomery reduction algorithm immediately. General principles of creating algorithms of 
Montgomery type are then suggested. In section 3, the framework described in
section 2 is used to analyze recent notable variants of Montgomery algorithm,
the validation of these methods can be performed easily. We conclude the paper in
section 4.

\section{The Chinese Remainder Theorem and Montgomery Algorithm}\label{sec2}
                                                                                  In this section, we described a framework to model Montgomery reduction algorithm and its variants.
As mentioned before, the famous Montgomery algorithm is usually thought as a generalization of
Hensel's odd division for $2$-adic numbers.
We observe that the expressions of Montgomery reduction algorithm and a
solution to the Chinese remainder theorem are identical in some sense, so they have a natural relation.
Further more, the recent notable variants of Montgomery reduction algorithm 
also have Chinese remainder theorem interpretation. So the CRT approach 
provides a unified, natural and transparent treatment to this family of algorithms.

The CRT is a well known method for solving a system of modular equation.
It is also known as  Sun Tzu Theorem as it was described in a very ancient Chinese book ``Sun Tzu Suan Jing''.
In  his  ``Mathematical Treatise in Nine Sections'' of 1247 \cite{qin1247},  Qin
described the Chinese remainder theorem with great detail and generality.

In \cite{qin1247}, Qin discussed the concept of `precise use'\footnote{In \cite{Libb} (page 331), `precise use' is termed as `ch\^eng-yung' to keep Chinese pronunciation.} , it can be summarized into
the following equality (\ref{eq:qin0}) for the case of two moduli. We remark that
although its proof is straightforward, the fact that Qin paid special attention to this expression
in his derivation of CRT is very interesting.

\begin{thm} (Qin's Identity)\label{qin:iden} Let $N, R>1$ be two coprime integers. Denote $N^{-1} = N^{-1}\pmod R, \ R^{-1}=R^{-1}\pmod N$,
then
	\begin{equation}\label{eq:qin0}
		R^{-1}R+N^{-1}N = 1+NR.
	\end{equation}
\end{thm}
\begin{proof}By definition, there exists a positive integer $\ell$ such that
	\[
	N^{-1}N = 1+\ell R.
	\]
	Since $1\le N^{-1}<R$, we see that $\ell <N$. Furthermore, $\ell R\equiv -1 \pmod N$, so $N-\ell=R^{-1}$. Therefore,
	\[
R^{-1}R +N^{-1}N=(N-\ell)R+N^{-1}N=1+NR.
	\]
\end{proof}

Multiplying a given an integer $T$ to the both side of (\ref{eq:qin0}), we see that 
\begin{equation}\label{eq:qin1}
TR^{-1}R+TN^{-1}N = T+TNR.
\end{equation}

We shall first have some discussion on the identity (\ref{eq:qin1}) here. This identity is essentially the formula of the Chinese Remainder Theorem for the system of congruence 
\[
\left\{ \begin{array}{l} T \equiv r_1 \pmod N\\
	T\equiv r_2\pmod R \end{array}\right.
\]
Note that $ TR^{-1}R+TN^{-1}N\equiv r_1R^{-1}R+r_2N^{-1}N \pmod {NR}$, so from
 (\ref{eq:qin1}) we see that 
 \[
 T\equiv r_1R^{-1}R+r_2N^{-1}N \pmod {NR}.
 \]
For the general case of multiple moduli, one can derive the CRT 
in a similar manner, see \cite{dlx}. 

Next, we explain that (\ref{eq:qin1}) actually derives  the Montgomery reduction algorithm. Writing $N'=R-N^{-1}$,  the identity (\ref{eq:qin1})  becomes
\begin{equation}\label{eq:qin2}
	TR^{-1}R=T+TN'N
\end{equation}
This means that $T+TN'N$ is divisible by $R$ and 
\[
	TR^{-1}=\frac{T+TN'N}R
\]
Recall that $m = TN' \pmod R$ in the  Montgomery reduction\footnote{In line 1 of the  Montgomery reduction, $m$ is calculated by $m = (T \pmod R)N' \pmod R$ to further reduce the time of multiplication. }, so
\begin{equation}\label{eq:qin3}
	TR^{-1}\equiv \frac{T+mN}R \pmod N.
\end{equation}
As it is  routine  to check that $\frac{T+mN}R=\frac{T}R+\frac{mN}R<2N$, (\ref{eq:qin3}) gives the celebrated Montgomery reduction algorithm.

This process is very natural in that $TR^{-1}R$ is a term in (\ref{eq:qin1})
and (\ref{eq:qin2}). Some general principles are suggested from the process. To this end, we first fix a notation {\bf mods} for signed remainder in this paper:  by $x ( \mbox{ mods } n)$, we mean the least absolute remainders of an integer $x$ dividing by a positive integer $n$ (which is an integer in the interval $[-\frac{n}2, \frac{n}2)$). 

We now set up a general framework for reduction algorithm of Montgomery type.
The task for Montgomery algorithm (and its variants) is to efficiently compute
$r= TR^{-1} \pmod N$, or $r=  TR^{-1} (\mbox{ mods } N)$, or
$r\equiv TR^{-1} \pmod N$ for integer $r$ in certain range (e.g., $a\le r<b$). By (\ref{eq:qin1}), that means that
\begin{equation}\label{eq:qin4}
r \equiv \frac{T-TN^{-1}N+TNR}R \pmod N  \equiv \frac{T-m N}R \pmod N .
\end{equation}
where $m$ is a suitable integer such that $m\equiv TN^{-1} \pmod R$. Some of the key points are
\begin{itemize}
	\item $T-m N$ is divisible by $R$.
	\item $m$ is to be selected so that $\frac{T-m N}R$ is small. Given that $|T| <RN$, this is true for $m= TN^{-1} \pmod R$, or $m=  TN^{-1} (\mbox{ mods } R)$.
	\item The desired return value is the number $\frac{T-m N}R$. This number may be computed more efficiently by utilizing the fact that $R$ is a power of two. 
\end{itemize}
It is noted that some of the algorithms of Montgomery type are required to calculate
$-TR^{-1} \pmod N$ or $- TR^{-1} (\mbox{ mods } N)$, so one just needs to take $-\frac{T-m N}R=\frac{m N - T}R$ as the return value.

\section{Analysis of Reduction Algorithms  of Montgomery Type}
The
number theory transform has been a very powerful tool for 
lattice-based encryption systems built on structured polynomial lattices. For this purpose, the computation of modular multiplications for small (e.g. word size) moduli has received recent attention and several notable reduction algorithms of Montgomery type have been proposed \cite{AMOT, Huang,Huang2,LyuSei,Plan,Seiler}. We find that the CRT approach can be used to model these algorithms
uniformly and to get natural derivation process.
The proofs of correctness become clear.  In particular, with this our approach, some problems of a modular reduction algorithm in \cite{Huang} are identified, and a counterexample is generated to show that the algorithm is incorrect as reported in an early version of the paper \cite{YJX}.

In our analysis, we will try to use the variable symbols  that are consistent with
the Montgomery algorithm.

\subsection{CRT Interpretation of Signed Montgomery Algorithm}

Recently, Seiler proposed a variation Montgomery reduction and used it to perform NTT for NTRU
 \cite{Seiler,LyuSei}. The modulus $R=2^{\ell}$
is restricted in the size of a word and
negative residue is allowed. The odd modulus $N$ satisfies $2N< R$.
The input $T$ with $|T|<\frac{NR}2$ is written as $T=a_1R+a_0$ with $0\le a_0<R$.
 The description of the Seiler algorithm is next.

\begin{algorithm}[htb]\small{
		\caption{Signed-Montgomery Reduction
			\label{alg-seiler}}
		\begin{algorithmic}[1]
			\Require{Integer $ -\frac{NR}2<T=a_1 R+a_0< \frac{NR}2, \ 0\le a_0<R$}
			\Ensure{ $ r\equiv TR^{-1} \pmod N$, $-N<r<N$.}
			\Function{SigRedc}{$T$}
			\Let{$m_0$}{$a_0N^{-1} ( \mbox{ mods } R)$ }
			\Let{$t$}{$\left\lfloor\frac{m_0N}{R}\right\rfloor$ }
			\Let{$r$}{$a_1-t$}
			\State \Return{$r$}
			\EndFunction
		\end{algorithmic}
	}
\end{algorithm}

Next we present a straightforward derivation of Algorithm \ref{alg-seiler}, together with a proof of its  correctness.

As discussed at the end of last section, the return value $r$ satisfies
\[
r\equiv TR^{-1} \pmod N \equiv \frac{T-mN}R \pmod N 
\]
with $m \equiv TN^{-1} ( \mbox{ mods } R)$. Since $T$ has a specific representation
of $T=a_1R+a_0$ with $0\le a_0<R$, we get 
\[
m  \equiv (a_1R+a_0)N^{-1} ( \mbox{ mods } R) \equiv a_0N^{-1} ( \mbox{ mods } R)=m_0.
\]
Now 
\[
\frac{T-mN}R \pmod N  \equiv \frac{a_1R+a_0-mN}R \pmod N =a_1- \frac{m_0N-a_0}R=r.
\]
This gives the desired Signed-Montgomery Reduction algorithm simply because
\begin{enumerate}
	\item $\left| a_1-\frac{mN-a_0}R \right|=\left| \frac{a_1R+a_0-mN}R \right|=\left| \frac{T-mN}R \right|
	<\frac{\frac{NR}2+\frac{R}2 N}R=N$, since $|m|\le \frac{R}2$.
	\item $\left\lfloor\frac{mN}{R}\right\rfloor = \frac{mN-a_0}R$, since $\frac{mN-a_0}R$ is an integer and $0\le \frac{a_0}R<1$.
\end{enumerate}

\subsection{CRT Interpretation of Plantard Reduction Algorithm and Its Extension}
In 2021,  Plantard designed a new modular reduction algorithm \cite{Plan}.
The method utilizes the special property of doing modular operation with in a word size. It
exhibits better performance for several cryptographic schemes. One of the nice features of the algorithm is to turn a multiplication of a $2n$-bit integer with an $n$-bit integer to a multiplication of two   $n$-bit integers.
To be more precise, let $\phi = \frac{1+\sqrt{5}}{2}$, $R=2^{2n}$. 
If the odd modulus $N$ satisfies $N<\frac{2^n}{\phi}$, then the following algorithm works.

\begin{algorithm}[htb]\small{
		\caption{Plantard Reduction Algorithm
			\label{alg-Plan}}
		\begin{algorithmic}[1]
			\Require{$N<\frac{2^n}{\phi},0 \leq T \leq N^2$, }
			\Ensure{$r = -TR^{-1}\pmod N$.}
			\Function{PRedc}{$T$}
			\Let{$r$}{$\left\lfloor\frac{\left(\left\lfloor\frac{ T N^{-1} \pmod R }{2^n}\right\rfloor+1\right) N}{2^n}\right\rfloor$ }
			\If {$(r = N)$}
			\State \Return{$0$}
			\EndIf
			\State \Return{$r$}
			\EndFunction
		\end{algorithmic}
	}
\end{algorithm}

To use the CRT model, we argue as follows.

From (\ref{eq:qin1}), we get
\[
	-TR^{-1}R = TN^{-1}N-T-TNR=\big(TN^{-1} \pmod R\big)N-T+kNR
\]
for some integer $k$. This means that $\big(TN^{-1} \pmod R\big)N-T$ is divisible by $R$ and
\[
-TR^{-1}\equiv \frac{\big(TN^{-1} \pmod R\big)N-T}R\pmod N
\]

We just need to verify that 
\begin{equation}\label{eq:Pl}
\frac{\big(TN^{-1} \pmod R\big)N-T}R=\left\lfloor\frac{\left(\left\lfloor\frac{ T N^{-1} \pmod R }{2^n}\right\rfloor+1\right) N}{2^n}\right\rfloor.
\end{equation}
Writing $m=TN^{-1}\pmod R$. Note that $R=2^{2n}$ and $0\le T<N^2<\frac{R}{\phi^2}$, so (\ref{eq:Pl}) is equivalent to 
\[
\frac{mN-T}{2^n}\le \bigg(\left\lfloor\frac{m }{2^n}\right\rfloor+1\bigg) N<\frac{mN+R-T}{2^n}.
\]
The first inequality is obvious. By the fact that $\frac{1}{\phi}=1-\frac{1}{\phi^2}$ we get   the second inequality:
\begin{eqnarray*}
	\frac{mN+R-T}{2^n} &\ge & \frac{mN+R-\frac{R}{\phi^2}}{2^n}=\frac{m }{2^n}N+2^n(1-\frac{1}{\phi^2})\\
	&=&\frac{m }{2^n}N+\frac{2^n}{\phi}>\bigg(\left\lfloor\frac{m }{2^n}\right\rfloor+1\bigg) N.
\end{eqnarray*}
Finally, since $0\le m<R, \ 0\le T <R$,  it is clear that $-1<\frac{mN-T}R<N$. Namely,
\[
0\le \frac{\big(TN^{-1} \pmod R\big)N-T}R<N.
\]

{\bf Remark}. It is interesting to note that we have acually shown that
$\left\lfloor\frac{\left(\left\lfloor\frac{ T N^{-1} \pmod R }{2^n}\right\rfloor+1\right)N}{2^n}\right\rfloor<N$, so the {\bf if} clause
in lines 3 and 4 of algorithm \ref{alg-Plan} can be removed. We remark that this has been pointed out in \cite{AMOT} through some calculation.

A further discussion on Plantard Reduction Algorithm is given in  \cite{AMOT}  where a signed version is proposed. The 
algorithm considers $R=2^{2n}$ and odd modulus $N$ such that $N<2^{n-1}$ and sets $\widetilde{N}^{-1} = N^{-1}  (\mbox{ mods } R),  \   \widetilde{R}^{-1} = R^{-1}  (\mbox{ mods } N)$. The goal is to
compute  $-T\widetilde{R}^{-1} (\mbox{ mods } N)$
for an input $T$ with $|T|<2^{2n-2}$.

\begin{algorithm}[htb]
		\caption{Signed Plantard Multiplication
			\label{alg-SigPlan}}
		\begin{algorithmic}[1]
			\Require{Integer $T$ with $ |T|\leq 2^{2n-2}$ }
			\Ensure{$r = -T\widetilde{R}^{-1}(\mbox{ mods } N)$.}
			\Function{SigPRedc1}{$T$}
			\Let{$r$}{$\left\lfloor\frac{\left\lfloor\frac{ T \widetilde{N}^{-1} (\mbox{ mods } R) }{2^n}\right\rceil N}{2^n}\right\rceil$ }
			\State \Return{$r$}
			\EndFunction
		\end{algorithmic}
\end{algorithm}

To explain this with CRT, we rewrite (\ref{eq:qin1})  as $T\widetilde{R}^{-1} R+T\widetilde{N}^{-1} N=T+kNR$ for some integer $k$. Therefore, $\big(T\widetilde{N}^{-1} (\mbox{ mods } R) \big)N-T$ is divisible by $R$ (with a smaller quotient) and 
\[
-T\widetilde{R}^{-1}(\mbox{ mods } N)\equiv \frac{\big(T\widetilde{N}^{-1} (\mbox{ mods } R) \big)N-T}{R}.
\]
Write $m=T\widetilde{N}^{-1} (\mbox{ mods } R)$,  we need to check that 
\[
\frac{mN-T}{R}=\left\lfloor\frac{\left\lfloor\frac{ m }{2^n}\right\rceil N}{2^n}\right\rceil
\]
Since $=\left\lfloor\frac{\left\lfloor\frac{ m }{2^n}\right\rceil N}{2^n}\right\rceil=\left\lfloor\frac{\left\lfloor\frac{ m }{2^n}+\frac{1}2\right\rfloor N}{2^n}+\frac{1}2\right\rfloor$, this is equivalent to checking
\[
\frac{mN-T}{2^n}\le \left\lfloor\frac{ m }{2^n}+\frac{1}2\right\rfloor N+2^{n-1}<\frac{mN+R-T}{2^n}
\]
To prove the first inequality, we use the facts of $N<2^{n-1}$ and $|T|<2^{2n-2}$:
\[
\frac{mN-T}{2^n}=\big(\frac{m}{2^n}-\frac{1}2\big)N+\frac{N}2-\frac{T}{2^n}\le \left\lfloor\frac{ m }{2^n}+\frac{1}2\right\rfloor N
+\frac{N}2+\frac{|T|}{2^n}< \left\lfloor\frac{ m }{2^n}+\frac{1}2\right\rfloor+2^{n-1}.
\]
To prove the second inequality, we note that $R-T-2^{n-1}N>2^{2n}-2^{2n-2}-2^{2n-2}=2^{2n-1}$, so  
\[
\left\lfloor\frac{ m }{2^n}+\frac{1}2\right\rfloor N+2^{n-1}\le
\big(\frac{ m }{2^n}+\frac{1}2\big) N+2^{n-1}<\frac{mN+R-T}{2^n}.
\]


\subsection{Discussion on Another Signed Version of Plantard Algorithm}
Observing that a modulus in many applications is usually much smaller than a word, Huang et al introduced another signed version of  Plantard algorithm 
to processes signed integers \cite{Huang}.  In the algorithm, $R$ is
$2^{2n}$, while
the odd modulus $N$ satisfies $N<2^{n- \alpha -1}$, where $\alpha \geq 0$ is an integer parameter. Write $\widetilde{N}^{-1} = N^{-1}  (\mbox{ mods } R),  \   \widetilde{R}^{-1} = R^{-1}  (\mbox{ mods } N)$, the goal is to
compute  $-T\widetilde{R}^{-1} (\mbox{ mods } N)$
for an input $T$ with $|T|<2^{2\alpha}N^2$.

\begin{algorithm}[htb]
		\caption{Signed Plantard Reduction Algorithm-(2)
			\label{alg-SigPlan2}}
		\begin{algorithmic}[1]
			\Require{Integer $T$ with $|T|\le 2^{2\alpha}N^2$, integer $\alpha\ge 0$}
			\Ensure{$r \equiv -T\widetilde{R}^{-1}(\mbox{ mods } N),-\frac{N}{2} < r <\frac{N}{2}$.}
			\Function{SigPRedc2}{$T,\alpha$}
			\Let{$r$}{$\left\lfloor\frac{\left(\left\lfloor\frac{ T \widetilde{N}^{-1} (\mbox{ mods } R) }{2^n}\right\rfloor +2^\alpha\right)N}{2^n}\right\rfloor$ }
			\State \Return{$r$}
			\EndFunction
		\end{algorithmic}
\end{algorithm}

We now use the CRT approach to analyze this method. Again, from (\ref{eq:qin1})  we get  $T\widetilde{R}^{-1} R+T\widetilde{N}^{-1} N=T+kNR$ for some integer $k$. and hence $\big(T\widetilde{N}^{-1} (\mbox{ mods } R) \big)N-T$ is divisible by $R$ (with a smaller quotient). Therefore 
\[
-T\widetilde{R}^{-1}(\mbox{ mods } N)\equiv \frac{\big(T\widetilde{N}^{-1} (\mbox{ mods } R) \big)N-T}{R}.
\]
Write $m=T\widetilde{N}^{-1} (\mbox{ mods } R)$,  if the algorithm were correct, we would have
\[
\frac{mN-T}{R}=\left\lfloor\frac{\left(\left\lfloor\frac{ m }{2^n}\right\rfloor +2^\alpha\right)N}{2^n}\right\rfloor.
\]
Namely, the following inequalities would be true
\[
\frac{mN-T}{2^n}\le \bigg(\left\lfloor\frac{m}{2^n}\right\rfloor +2^\alpha \bigg)N<\frac{mN+R-T}{2^n}
\]
The second inequality is always true by the facts that $R>2^{n+\alpha+1}N$ and $|T|\le 2^{n+\alpha-1}N$:
\[
\frac{mN+R-T}{2^n}> \left\lfloor\frac{m}{2^n}N\right\rfloor+\frac{2^{n+\alpha+1}N - 2^{n+\alpha-1}N }{2^n}> \left\lfloor\frac{m}{2^n}\right\rfloor N+2^\alpha N.
\]
The first one holds for $\alpha \ge 1$ since
\[
\frac{mN-T}{2^n}< \bigg(\left\lfloor\frac{m}{2^n}\right\rfloor +1 \bigg)N+\frac{|T|}{2^n}< \bigg(\left\lfloor\frac{m}{2^n}\right\rfloor +1 \bigg)N+
2^{\alpha-1}N\le \bigg(\left\lfloor\frac{m}{2^n}\right\rfloor +2^{\alpha} \bigg)N
\]
For the case of $\alpha = 0$, however, $\frac{mN-T}{2^n}\le  \bigg(\left\lfloor\frac{m}{2^n}\right\rfloor +1 \bigg)N$ is not true in general. We construct a counter example.
\begin{example} Let $n = 6, N=31, \alpha=0$. So $R=2^{2n}=4096$ and $N$ satisfies $N<2^{n-\alpha-1}=32$.
	
It is calculated that $\widetilde{N}^{-1} = N^{-1}  (\mbox{ mods } R)= -1057$> Now we take inputs $T=-95$, then $m=-1985=(-95)(-1057)  (\mbox{ mods } R)$ and
\[
\frac{mN-T}{2^n}= -15\times 64 = -960.
\]
But 
\[
\bigg(\left\lfloor\frac{m}{2^n}\right\rfloor +1 \bigg)N = -961.
\]

In this example, the return value of the algorithm is
\[
\left\lfloor\frac{\left(\left\lfloor\frac{ T \widetilde{N}^{-1} (\mbox{ mods } R) }{2^n}\right\rfloor +2^0\right)N}{2^n}\right\rfloor=-16,
\]
which is not congruent to $-T\widetilde{R}^{-1}$ modulo $N$, and not even in the required interval
$[-15,15]$. So algorithm \ref{alg-SigPlan2} is incorrect.
\end{example}

{\bf Remarks}
\begin{enumerate}
\item The original statement of algorithm \ref{alg-SigPlan2} \cite{Huang} contains some incorrect use of notations as well. These are also cleared in the our early version \cite{YJX} of this paper, so this part is not touched on here. 
\item After learning \cite{YJX}, the authors of \cite{Huang} changed their
parameter to $\alpha\ge 1$, this is also reflected in their more recent paper\cite{Huang2}. We see that the CRT approach is to design an efficient computation of $\frac{mN-T}{R}$, the identification of an error for $\alpha=0$
and correctness for $\alpha>0$ become very transparent, as done in the above discussion.
\end{enumerate}

\section{Conclusion}
In this paper, we setup a CRT framework to treat Montgomery reduction and its variants.
Under this approach, the derivation of these algorithms can be revealed more transparently, their
proof of correctness can be processed naturally.
Using this approach,
some problems of the modular reduction algorithm in
the literature are identified, a counterexample is generated to show that the algorithm is incorrect.


\end{document}